The cs4rep Class for L^{A}TEX 2ε

Martin Reddy

April 25, 1996

1 Introduction

This document describes the cs4rep class which provides a document style for presenting 4th year undergraduate project reports in the department of Computer Science. To use this class, all you need do is begin your L^{A}TEX document with the declaration:

     \documentclass[<options>]{cs4rep}

The default behaviour of this formatting class is to produce documents which are typeset as 12 point single-spaced text on double-sided A4 pages (with all new chapters appearing on the first clear right-hand page). Note that double-sided reports can be printed using the -Zduplex option to lpr on laser printers that support duplex printing.

The reader’s attention is also drawn to the csthesis style file which provides a generic document style for submitting theses in the Computer Science department which are in line with the official University regulations (see “texdoc csthesis” for further details).

2 Class Options

Class options are specified in the the square-bracketed section of the \documentclass command. The cs4rep class provides one class option, and allows only a couple of the standard L^{A}TEX 2ε report class options to be passed (e.g. leqno and fleqn). The following class option is currently supported by cs4rep:

parskip : alters the paragraph formatting so that each paragraph is separated by a vertical space, and there is no indentation at the start of each paragraph. (The list environments are also updated to reflect this change.)

3 The Title Page

The cs4rep class produces a title page which is formatted so that it may be bound using the standard departmental (orange) report covers. When this is the case, the project’s title, author and submission date will be visible, along with the header “4th Year Project Report”.

The following short example illustrates how to produce a title page using the cs4rep class. Note that the abstract is displayed at the bottom of the title page and so must be defined before executing a \maketitle command.

  \title{<report_title>}
  \author{<student's_name>}
  \date{<date_of_submission>}
  \abstract{<report_abstract>}
  \maketitle

Each of the above four fields can include a \thanks{<text>} attribution to provide any additional comments (these will be formatted as a footnote at the bottom of the title page).

4 Other Features

The cs4rep class does not allow \listoffigures or \listoftables. These are deemed unnecessary for CS4 reports.